Circular No. 5671

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SWIFT-TUTTLE (1992t) Further to IAUC 5670, Waddington, considering several assumed val-

ues of A_{2} between –0.0075 and +0.0075, has independently shown that A_{2} = −0.0025 yields perihelion dates T = +188 Aug. 17 and –68 July 15; although the computation for –68 is heavily influenced by a close approach to the earth in +188, this solution is significant in that it yields no other approaches within 0.5 AU of the earth between +188 and 1737. The general stability of Waddington’s result is confirmed by G. V. Williams (Harvard-Smithsonian Center for Astrophysics = CfA), who finds that the actual orbit and similar nongravitational parameters on IAUC 5670 yield T = +188 July 20 and an earth approach to 0.23 AU 12 days later; with an arbitrary change of up to –20 days in T a closer approach and a representation of the –68 observations becomes possible. J. E. Chambers (Manchester University and CfA) finds the comet’s long-term motion to be dominated by the 1:11 mean-motion resonance with Jupiter, in spite of the nongravitational effects and occasional encounters with the earth and Saturn. Although post-perihelion observations in 1993 and subsequently should be examined for a possible repetition of the systematic residuals of 1862 Oct., the proposed identity with the +188 comet suggests that the next two returns of P/Swift-Tuttle will likely be within a day of T = 2126 July 11 and 2261 Aug. 12 and that there is very little possibility of the comet’s collision with the earth during the next millennium.

1992 /93 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 4 19^{h}09.^{m}67 + 0^{o}51.^{′}3 1.428 0.969 42.^{o}6 43.^{o}5 5.1
6 19 15.08 − 1 08.7 1.462 0.964 40.9 42.0 5.1
8 19 20.25 − 3 03.2 1.496 0.961 39.2 40.4 5.1
10 19 25.20 − 4 52.6 1.530 0.959 37.5 38.7 5.2
12 19 29.95 − 6 37.2 1.565 0.958 35.8 36.9 5.2
14 19 34.51 − 8 17.1 1.600 0.959 34.1 35.1 5.2
16 19 38.91 − 9 52.6 1.634 0.960 32.4 33.3 5.3
18 19 43.17 −11 24.1 1.668 0.963 30.7 31.4 5.4
Feb. 12 21 24.70 −39 17.4 2.221 1.400 25.9 17.9 8.4
17 21 34.65 −41 13.0 2.229 1.457 29.8 19.7 8.7
22 21 45.06 −43 08.6 2.233 1.515 33.9 21.3 9.0
27 21 56.01 −45 04.7 2.233 1.574 37.9 22.8 9.2
Mar. 4 22 07.58 −47 02.1 2.229 1.634 42.0 24.0 9.4

1992 December 5 Brian G. Marsden